OpenStudy (anonymous):

Simplify the expression below:

OpenStudy (anonymous):

\[\frac{ \sqrt{7} }{ \sqrt[4]{7} }\]

OpenStudy (anonymous):

\[7^\frac{- 3 }{ 4 }\] \[7^\frac{1}{ 4 }\] \[7^\frac{3 }{ 4 }\] 7

OpenStudy (anonymous):

(Second comment is answer choices)

OpenStudy (anonymous):

I'm pretty sure it's B (7^1/4) but I'm not positive. Or if you need it as a decimal it's something like 1.6265